### v3

@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive advancements across a wide range of natural language processing tasks, including syntactic parsing (Ma et al., 2024), complex scientific reasoning (Wang et al., 2023), and commonsense knowledge answering (Zhao et al., 2023). As the model size and training data scale up, LLMs have demonstrated the ability to surpass human-level accuracy on certain benchmarks (Srivastava et al., 2022), highlighting their emerging capacity for sophisticated reasoning and problem-solving.
To better enhance LLM reasoning capabilities–and to push their performance closer to, or even beyond, human-level reasoning–numerous prompting-based strategies have been proposed. Chain-of-Thought (CoT) prompting encourages explicit decomposition of complex problems into intermediate steps (Wei et al., 2022; Zhou et al., 2022), while Tree-of-Thought (ToT) generalizes this idea by exploring multiple branching reasoning paths (Yao et al., 2023). Sampling-based approaches like Best-of-N improve robustness by selecting the most coherent reasoning path from diverse candidates (Ji et al., 2023), and automatic prompt optimization techniques aim to systematically discover prompts that better facilitate multi-step reasoning (Zhang et al., 2022; Shum et al., 2023). If samples of the same type of question are provided, In-Context Learning (ICL) (Brown et al., 2020) also prompts LLM with few-shot examples with advanced performance.
Despite these advances, LLM reasoning remains highly configuration-sensitive: as Figure 1 shows, GPT-4o’s accuracy on the metaphor expression classification task (Tong et al., 2024) swings wildly under different reasoning configurations. While divergent reasoning prompts and fewer reasoning steps could greatly improve performance, temperature as 1 instead drown out useful reasoning with noise, negating any benefit from the added randomness. However, previous methods have not targeted tuning on these parameters. CoT (Wei et al., 2022; Zhou et al., 2022) and ToT (Yao et al., 2023) apply fixed reasoning structures that fail to generalize to creative or subjective tasks (e.g. spatial planning (Stechly et al., 2024)). Best-of-N (Ji et al., 2023) rely on unguided generation, suffering from a “garbage in, garbage out” effect. Automatic prompt optimization (Zhang et al., 2022; Shum et al., 2023) focuses on static templates and overlooks crucial hyperparameters like temperature, failing to adjust reasoning strategies. While ICL (Brown et al., 2020) extracts some cues from input questions, it remains brittle under context perturbations (Mueller et al., 2023), and its reliance on implicit pattern matching has been shown to be less effective than direct structured reasoning (Tang et al., 2023). These limitations call a need for an adaptive prompting configuration strategy for LLMs to handle various sophisticated reasoning.
Figure 1: Performance of different CoT settings on the metaphor dataset (Tong et al., 2024). The default temperature is 0.1 if not specified.
@@
We introduce AdaReasoner [^1] Implementation code is avaliable at: https://github.com/XiangqiWang77/officialadareasoner, an LLM-agnostic plugin designed to automate adaptive reasoning configurations for tasks requiring diverse types of thinking. When integrated with an LLM, AdaReasoner is trained using a reinforcement learning (RL) framework. In this setup, AdaReasoner acts as a decision-making agent, where the state is defined by the current task presented to the LLM, reflecting the nature of the reasoning required (e.g., logical, creative, or subjective). The action corresponds to selecting a configuration from an action space composed of three key hyperparameters: (i) the reasoning instruction formats, (ii) the generation temperature, and (iii) the number of reasoning steps. To enable AdaReasoner to learn the most effective configuration policy, a pretrained reward model is employed to evaluate the effectiveness of the reasoning configuration. This model provides feedback to guide the agent’s learning, enabling it to efficiently acquire effective configurations with only limited guidance (i.e., few-shot learning). To facilitate exploration and improve generalization, we employ a Boltzmann exploration mechanism, enabling the agent to explore and optimize configurations more effectively during training. Once trained, AdaReasoner is used as a plug-in to the LLM, providing adaptive reasoning configurations that allow the model to adjust its reasoning approach based on the task at hand.
@@
  • We introduce AdaReasoner, an LLM-agnostic plugin that automates adaptive reasoning configurations for tasks requiring diverse types of thinking. s a reinforcement learning framework with a factorized action space. Its training is data-efficient yet scalable, requiring only a small number of samples for each task aided by the use of the Boltzmann exploration mechanism.
@@ 2 Related Work of Reasoning in LLMs
The pursuit of enhanced reasoning capabilities in LLMs has spurred diverse research trajectories, beginning with foundational techniques like Chain-of-Thought (CoT) prompting (Wei et al., 2022). CoT enables LLMs to articulate intermediate steps, significantly improving performance on complex tasks. However, its efficacy can be hampered by sensitivity to prompt formulation (Sprague et al., 2024; Puerto et al., 2024) and limitations in subjective or creative domains (Chochlakis et al., 2024; Xu et al., 2024), sometimes even degrading performance where brevity is key (Liu et al., 2024). To mitigate these issues and reduce manual effort, innovations such as Automatic CoT (Auto-CoT) (Zhang et al., 2022; Shum et al., 2023) emerged, automating the generation of effective reasoning exemplars. Further advancements include structured reasoning frameworks like Tree-of-Thoughts (ToT) (Yao et al., 2023) and Graph-of-Thoughts (GoT) (Besta et al., 2024), which allow models to explore and evaluate multiple reasoning pathways, alongside methods like CoT-influx (Huang et al., 2023) that optimize few-shot CoT contexts.
@@
To bolster the robustness and reliability of LLM reasoning, researchers have explored self-correction and learning-based paradigms. Self-consistency techniques (Wang et al., 2022), often realized through Best-of-N sampling, leverage the generation of multiple diverse reasoning paths and subsequent aggregation (e.g., via majority voting) to improve answer accuracy. Complementary to this, self-reflection mechanisms, as seen in Self-Refine (Madaan et al., 2023) and Reflexion (Shinn et al., 2023), empower LLMs to iteratively critique and enhance their own outputs, akin to human error correction, with some approaches fine-tuning with divergent CoT to specifically boost these capabilities (Puerto et al., 2024). Reinforcement Learning (RL) has also become a cornerstone for optimizing reasoning, from general alignment via RLHF (Ouyang et al., 2022) to specialized reward models that guide the LLM towards more accurate and effective thought processes (Jaech et al., 2024). Models like DeepSeek-R1 (Guo et al., 2025) exemplify LLMs fine-tuned with RL to excel at intricate reasoning, sometimes learning to control their own reasoning flow through meta-actions.
@@
The nuanced control of generation parameters and adaptive hyperparameter tuning represent another critical frontier. The stochastic decoding settings, such as temperature, significantly affect output diversity and, consequently, reasoning quality and creativity (Renze, 2024a). Higher diversity can fuel methods like self-consistency but requires careful management to maintain coherence. Recent work has thus focused on automated optimization of prompt configurations, decoding parameters, and even enabling LLMs to self-regulate their generation strategies, as demonstrated by Hyperparameter-Aware Generation (HAG) (Wang et al., 2024). Our AdaReasoner contributes to this line of research by introducing an adaptive framework that explicitly manages a toolbox of reasoning hyperparameters, including the reasoning method prompt, temperature, and number of reasoning steps, using an RL-trained agent to dynamically tailor the reasoning process to individual inputs, coupled with self-reflection and a robust selection mechanism for enhanced flexibility.
@@ 3 AdaReasoner
Motivation. Even though CoT and similar LLM reasoning methods have been studied as generally efficient and helpful, they still cannot achieve ideal performance across all types of questions. For example, tasks like joke generation or metaphor interpretation often require divergent and creative reasoning chain  (Zhong et al., 2024). For more complex reasoning tasks, stronger and more explicit reasoning instructions would be beneficial (Lin et al., 2024). Thus, adapting LLM configurations tailored for specific tasks is crucial for achieving better overall performance. As illustrated in Figure 2, we design AdaReasoner to adapt reasoning configurations by taking actions as a combination of different hyperparameters for LLMs. The inference/evaluation process is illustrated by the black arrows, while the training flow is depicted by the cyan arrows.
@@ 3.1 Hyperparameter Configuration (Action)
As mentioned earlier, we consider three hyperparameters in the reasoning configuration: 1) the generation temperature $a_{t}$; 2) the format of reasoning instructions $a_{p}$; and 3) the number of reasoning steps in CoT $a_{s}$, for several reasons. First, they have substantial impacts on the reasoning performance. Previous studies have revealed that the generation temperature modulates the diversity of model outputs, often yielding markedly different responses when varied (Renze, 2024b). The number of reasoning steps reflects the depth and thoroughness of the inference process and it thus could influence the reasoning accuracy (Dutta et al., 2024; Jin et al., 2024). The format of reasoning instructions, such as backward reasoning and step-by-step deduction, also plays a crucial role in guiding the model’s reasoning process (Almeida et al., 2024; Wang, 2025). Second, the settings of these three hyperparameters are adaptable for both proprietary and open-weight LLMs, with enhancement of adareasoner’s versatility. Third, we are aware of other hyperparameters that may also impact reasoning, such as the $p$ in top-$p$ sampling during generation and the random seed. However, we exclude $p$ because tuning top-$p$ alongside temperature is not recommended together with temperature aut . Additionally, our empirical evaluation found that varying the random seed could not be beneficial for improving LLMs’ reasoning performance (as shown in Section 4.3).
@@
To ensure practical feasibility, these configuration actions are discretized with a finite set of options. Specifically, the number of reasoning steps is bounded to avoid extreme values, defined as $\mathcal{A}_{s}$ as integers set, and temperature is discretized as set $\mathcal{A}_{t}$. The options for reasoning instructions, denoted as $\mathcal{A}_{p}$, are constructed based on a compositional design grounded in structure-mapping theory from cognitive psychology (Gentner, 1983), which models human reasoning by composing a core reasoning structure with contextual modifications. Accordingly, each reasoning instruction is factorized into two components: a base component, which specifies the overall cognitive strategy (e.g., creative thinking, analogical mapping, self-audit (Byrne et al., 2019)), and a variation, which modulates the emphasis on specific parts of the question or modifies the reasoning surface form. For example, a base “Apply creative reasoning to unearth unconventional insights and challenge standard assumptions” could be combined with a variation “Use simple, straightforward language to guarantee clarity and accessibility” for guiding the reasoning of divergent thinking types of problems. The same base, when combined with a variation Validate conclusions by aligning them with established principles or empirical data, such instruction is useful for critical thinking types of reasoning problems. Detailed of the base and variation components and their instantiation are provided in Appendix C. The reasoning instruction action space, $\mathcal{A}_{p}$, is composed of pairs in the form of {base, variation}. Each action $a_{p}$ corresponds to one of the possible combinations of a base and its associated variation.
@@ 3.2 Design and Training of AdaReasoner
Neural Architecture of AdaReasoner. As shown in Figure 2, the input query question, after embedding, undergoes three action selections before being sent to the LLMs for reasoning to generate the answer. While the embedding is performed (e.g. by pre-trained BERT model (Wolf et al., 2020)), the trainable neural network parameters of AdaReason consist of three parallel channels, each corresponding to one action, and one shared common layer as in Figure 2. The workflow is as follows: let $Embed(q)$ be the embedding of the input question $q$. It is first passed through the common layer to obtain $h=f_{\theta_{c}}(Embed(q))$, where $\theta_{c}$ are the parameters of the common layer (e.g., a fully connected MLP), and $h$ captures the features necessary to determine the actions.
@@
Exploration Strategy. By formulating the configuration selection for each question as a multi-armed bandit (MAB) problem, we aim to design an effective exploration strategy under the few-shot training setting. However, since the reward is derived indirectly from LLM outputs and the process is not an online learning scenario, standard MAB strategies such as Upper Confidence Bound (UCB) (Sutton and Barto, 2018) become impractical. Moreover, evaluating all configurations for each context $q$ is computationally infeasible, especially given the noisy and implicit reward landscape induced by LLM responses. Therefore, it is crucial to explore broadly across the configuration space while still prioritizing high-reward actions, and Boltzmann exploration offers an effective solution (Pan et al., 2019), as it allows the agent to probabilistically select actions based on their estimated rewards. Specifically, for each action ($a_{t}$, $a_{s}$ or $a_{p}$), we estimate the selection probability for its all possible values (in $\mathcal{A}_{t}$, $\mathcal{A}_{s}$ or $\mathcal{A}_{p}$),
@@
where $Q(a_{i})$ is the logit score in the output layer of one policy network $f_{\theta}$ for action $a_{i}$. The temperature $\tau$ in Boltzmann exploration controls the exploration-exploitation trade-off: higher $\tau$ promotes exploration, lower $\tau$ favors exploitation. We anneal $\tau$ exponentially as $\tau_{t}=\tau_{0}\cdot\alpha^{t},t\leq T$, allowing the policy to gradually shift from broad exploration to reliable configuration selection and refined optimization (Kirkpatrick et al., 1983).
Reward Signal. Similar to previous work (Kwon et al., 2023; Ma et al., 2023; Rocamonde et al., 2023) using pre-trained language model as reward on light-weight RL model, we employ a language judgement model (ours is DeBERTa-based) as reward model (OpenAssistant, 2023) to provide feedback on the selected actions. Concretely, for the resulting generated answer $\Phi(q|a)$, it is presented to the reward model alongside the original question $q$ and reference answer $R$ in the form of the prompt “For $q$, the generated answer $\Phi(q|a)$ matches the ground truth $R$ and is correct”. The reward is computed from the model’s logits, providing a scalar score that enables fine-grained, differentiable supervision over diverse reasoning trajectories.
With the reward $r$, the AdaReasoner is optimized using the gradient descent (REINFORCE) algorithm (Silver et al., 2014), where the overall policy $\Pi_{\Theta}(a\mid q,\Phi)$ is factorized into three heads with a shared feature extractor $f_{\theta_{c}}$, and $\Theta=\{\theta_{c},\theta_{p},\theta_{t},\theta_{s}\}$ denotes the complete set of trainable parameters. For each head $j\in\{p,t,s\}$, we define the head-specific loss as $\mathcal{L}_{j}=-r\log\Pi_{\theta_{j}}(a\mid q,\Phi)$, resulting in a total loss $\mathcal{L}=\sum_{j\in\{p,t,s\}}\mathcal{L}_{j}$. The gradients are then computed via the chain rule, where the shared-layer gradient is aggregated as $\nabla_{\theta_{c}}\mathcal{L}=\sum_{j\in\{p,t,s\}}\nabla_{\theta_{c}}\mathcal{L}_{j}$, and used for updating
@@
This training scheme ensures that each sub-policy is guided by its own loss while the shared feature extractor $f_{\theta_{c}}$ is jointly optimized by all heads, thereby promoting coherence across the three action dimensions and preventing convergence to conflicting optima, in line with findings from multi-task learning (Ruder, 2017). Further training details are described in Algorithm 1.

### v4

@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive advancements across a wide range of natural language processing tasks, including syntactic parsing [26], complex scientific reasoning [52], and commonsense knowledge answering [59]. As the model size and training data scale up, LLMs have demonstrated the ability to surpass human-level accuracy on certain benchmarks [45], highlighting their emerging capacity for sophisticated reasoning and problem-solving.
To better enhance LLM reasoning capabilities–and to push their performance closer to, or even beyond, human-level reasoning–numerous prompting-based strategies have been proposed. Chain-of-Thought (CoT) prompting encourages explicit decomposition of complex problems into intermediate steps [54, 62], while Tree-of-Thought (ToT) generalizes this idea by exploring multiple branching reasoning paths [57]. Sampling-based approaches like Best-of-N improve robustness by selecting the most coherent reasoning path from diverse candidates [16], and automatic prompt optimization techniques aim to systematically discover prompts that better facilitate multi-step reasoning [58, 42]. If samples of the same type of question are provided, In-Context Learning (ICL) [5] also prompts LLM with few-shot examples with advanced performance.
Despite these advances, LLM reasoning remains highly configuration‐sensitive: as Figure 1 shows, GPT‑4o’s accuracy on the metaphor expression classification task [49] swings wildly under different reasoning configurations. While divergent reasoning prompts and fewer reasoning steps could greatly improve performance, temperature as 1 instead drown out useful reasoning with noise, negating any benefit from the added randomness. However, previous methods have not targeted tuning on these parameters. CoT [54, 62] and ToT [57] apply fixed reasoning structures that fail to generalize to creative or subjective tasks (e.g. spatial planning [46]). Best-of-N [16] rely on unguided generation, suffering from a “garbage in, garbage out” effect. Automatic prompt optimization [58, 42] focuses on static templates and overlooks crucial hyperparameters like temperature, failing to adjust reasoning strategies. While ICL [5] extracts some cues from input questions, it remains brittle under context perturbations [28], and its reliance on implicit pattern matching has been shown to be less effective than direct structured reasoning [48]. These limitations call a need for an adaptive prompting configuration strategy for LLMs to handle various sophisticated reasoning.
Figure 1: Performance of different CoT settings on the metaphor dataset [49]. The default temperature is 0.1 if not specified.
@@
We introduce AdaReasoner, an LLM-agnostic plugin designed to automate adaptive reasoning configurations for tasks requiring diverse types of thinking. When integrated with an LLM, AdaReasoner is trained using a reinforcement learning (RL) framework. In this setup, AdaReasoner acts as a decision-making agent, where the state is defined by the current task presented to the LLM, reflecting the nature of the reasoning required (e.g., logical, creative, or subjective). The action corresponds to selecting a configuration from an action space composed of three key hyperparameters: (i) the reasoning instruction formats, (ii) the generation temperature, and (iii) the number of reasoning steps. To enable AdaReasoner to learn the most effective configuration policy, a pretrained reward model is employed to evaluate the effectiveness of the reasoning configuration. This model provides feedback to guide the agent’s learning, enabling it to efficiently acquire effective configurations with only limited guidance (i.e., few-shot learning). To facilitate exploration and improve generalization, we employ a Boltzmann exploration mechanism, enabling the agent to explore and optimize configurations more effectively during training. Once trained, AdaReasoner is used as a plug-in to the LLM, providing adaptive reasoning configurations that allow the model to adjust its reasoning approach based on the task at hand.
@@
  • We introduce AdaReasoner, an LLM-agnostic plugin that automates adaptive reasoning configurations for tasks requiring diverse types of thinking.
  • AdaReasoner leverages a reinforcement learning framework with a factorized action space. Its training is data-efficient yet scalable, requiring only a small number of samples for each task aided by the use of the Boltzmann exploration mechanism.
@@ 2 Related Work of Reasoning in LLMs
The pursuit of enhanced reasoning capabilities in LLMs has spurred diverse research trajectories, beginning with foundational techniques like Chain-of-Thought (CoT) prompting [54]. CoT enables LLMs to articulate intermediate steps, significantly improving performance on complex tasks. However, its efficacy can be hampered by sensitivity to prompt formulation [44, 33] and limitations in subjective or creative domains [7, 56], sometimes even degrading performance where brevity is key [24]. To mitigate these issues and reduce manual effort, innovations such as Automatic CoT (Auto-CoT) [58, 42] emerged, automating the generation of effective reasoning exemplars. Further advancements include structured reasoning frameworks like Tree-of-Thoughts (ToT) [57] and Graph-of-Thoughts (GoT) [4], which allow models to explore and evaluate multiple reasoning pathways, alongside methods like CoT-influx [13] that optimize few-shot CoT contexts.
@@
To bolster the robustness and reliability of LLM reasoning, researchers have explored self-correction and learning-based paradigms. Self-consistency techniques [53], often realized through Best-of-N sampling, leverage the generation of multiple diverse reasoning paths and subsequent aggregation (e.g., via majority voting) to improve answer accuracy. Complementary to this, self-reflection mechanisms, as seen in Self-Refine [27] and Reflexion [41], empower LLMs to iteratively critique and enhance their own outputs, akin to human error correction, with some approaches fine-tuning with divergent CoT to specifically boost these capabilities [33]. Reinforcement Learning (RL) has also become a cornerstone for optimizing reasoning, from general alignment via RLHF [31] to specialized reward models that guide the LLM towards more accurate and effective thought processes [15]. Models like DeepSeek-R1 [11] exemplify LLMs fine-tuned with RL to excel at intricate reasoning, sometimes learning to control their own reasoning flow through meta-actions.
@@
The nuanced control of generation parameters and adaptive hyperparameter tuning represent another critical frontier. The stochastic decoding settings, such as temperature, significantly affect output diversity and, consequently, reasoning quality and creativity [35]. Higher diversity can fuel methods like self-consistency but requires careful management to maintain coherence. Recent work has thus focused on automated optimization of prompt configurations, decoding parameters, and even enabling LLMs to self-regulate their generation strategies, as demonstrated by Hyperparameter-Aware Generation (HAG) [51]. Our AdaReasoner contributes to this line of research by introducing an adaptive framework that explicitly manages a toolbox of reasoning hyperparameters, including the reasoning method prompt, temperature, and number of reasoning steps, using an RL-trained agent to dynamically tailor the reasoning process to individual inputs, coupled with self-reflection and a robust selection mechanism for enhanced flexibility.
@@ 3 AdaReasoner
Motivation. Even though CoT and similar LLM reasoning methods have been studied as generally efficient and helpful, they still cannot achieve ideal performance across all types of questions. For example, tasks like joke generation or metaphor interpretation often require divergent and creative reasoning chain  [61]. For more complex reasoning tasks, stronger and more explicit reasoning instructions would be beneficial [22]. Thus, adapting LLM configurations tailored for specific tasks is crucial for achieving better overall performance. As illustrated in Figure 2, we design AdaReasoner to adapt reasoning configurations by taking actions as a combination of different hyperparameters for LLMs. The inference/evaluation process is illustrated by the black arrows, while the training flow is depicted by the cyan arrows.
@@ 3.1 Hyperparameter Configuration (Action)
As mentioned earlier, we consider three hyperparameters in the reasoning configuration: 1) the generation temperature $a_{t}$; 2) the format of reasoning instructions $a_{p}$; and 3) the number of reasoning steps in CoT $a_{s}$, for several reasons. First, they have substantial impacts on the reasoning performance. Previous studies have revealed that the generation temperature modulates the diversity of model outputs, often yielding markedly different responses when varied [36]. The number of reasoning steps reflects the depth and thoroughness of the inference process and it thus could influence the reasoning accuracy [9, 17]. The format of reasoning instructions, such as backward reasoning and step-by-step deduction, also plays a crucial role in guiding the model’s reasoning process [2, 50]. Second, the settings of these three hyperparameters are adaptable for both proprietary and open-weight LLMs, with enhancement of adareasoner’s versatility. Third, we are aware of other hyperparameters that may also impact reasoning, such as the $p$ in top-$p$ sampling during generation and the random seed. However, we exclude $p$ because tuning top-$p$ alongside temperature is not recommended together with temperature [1]. Additionally, our empirical evaluation found that varying the random seed could not be beneficial for improving LLMs’ reasoning performance (as shown in Section 4.3).
@@
To ensure practical feasibility, these configuration actions are discretized with a finite set of options. Specifically, the number of reasoning steps is bounded to avoid extreme values, defined as $\mathcal{A}_{s}$ as integers set, and temperature is discretized as set $\mathcal{A}_{t}$. The options for reasoning instructions, denoted as $\mathcal{A}_{p}$, are constructed based on a compositional design grounded in structure-mapping theory from cognitive psychology [10], which models human reasoning by composing a core reasoning structure with contextual modifications. Accordingly, each reasoning instruction is factorized into two components: a base component, which specifies the overall cognitive strategy (e.g., creative thinking, analogical mapping, self-audit [6]), and a variation, which modulates the emphasis on specific parts of the question or modifies the reasoning surface form. For example, a base “Apply creative reasoning to unearth unconventional insights and challenge standard assumptions” could be combined with a variation “Use simple, straightforward language to guarantee clarity and accessibility” for guiding the reasoning of divergent thinking types of problems. The same base, when combined with a variation Validate conclusions by aligning them with established principles or empirical data, such instruction is useful for critical thinking types of reasoning problems. Detailed of the base and variation components and their instantiation are provided in Appendix C. The reasoning instruction action space, $\mathcal{A}_{p}$, is composed of pairs in the form of {base, variation}. Each action $a_{p}$ corresponds to one of the possible combinations of a base and its associated variation.
@@ 3.2 Design and Training of AdaReasoner
Neural Architecture of AdaReasoner. As shown in Figure 2, the input query question, after embedding, undergoes three action selections before being sent to the LLMs for reasoning to generate the answer. While the embedding is performed (e.g. by pre-trained BERT model [55]), the trainable neural network parameters of AdaReason consist of three parallel channels, each corresponding to one action, and one shared common layer as in Figure 2. The workflow is as follows: let $Embed(q)$ be the embedding of the input question $q$. It is first passed through the common layer to obtain $h=f_{\theta_{c}}(Embed(q))$, where $\theta_{c}$ are the parameters of the common layer (e.g., a fully connected MLP), and $h$ captures the features necessary to determine the actions.
@@
Exploration Strategy. By formulating the configuration selection for each question as a multi-armed bandit (MAB) problem, we aim to design an effective exploration strategy under the few-shot training setting. However, since the reward is derived indirectly from LLM outputs and the process is not an online learning scenario, standard MAB strategies such as Upper Confidence Bound (UCB) [47] become impractical. Moreover, evaluating all configurations for each context $q$ is computationally infeasible, especially given the noisy and implicit reward landscape induced by LLM responses. Therefore, it is crucial to explore broadly across the configuration space while still prioritizing high-reward actions, and Boltzmann exploration offers an effective solution [32], as it allows the agent to probabilistically select actions based on their estimated rewards. Specifically, for each action ($a_{t}$, $a_{s}$ or $a_{p}$), we estimate the selection probability for its all possible values (in $\mathcal{A}_{t}$, $\mathcal{A}_{s}$ or $\mathcal{A}_{p}$),
@@
where $Q(a_{i})$ is the logit score in the output layer of one policy network $f_{\theta}$ for action $a_{i}$. The temperature $\tau$ in Boltzmann exploration controls the exploration-exploitation trade-off: higher $\tau$ promotes exploration, lower $\tau$ favors exploitation. We anneal $\tau$ exponentially as $\tau_{t}=\tau_{0}\cdot\alpha^{t},t\leq T$, allowing the policy to gradually shift from broad exploration to reliable configuration selection and refined optimization [19].
Reward Signal. Similar to previous work [20, 25, 37] using pre-trained language model as reward on light-weight RL model, we employ a language judgement model (ours is DeBERTa-based) as reward model [30] to provide feedback on the selected actions. Concretely, for the resulting generated answer $\Phi(q|a)$, it is presented to the reward model alongside the original question $q$ and reference answer $R$ in the form of the prompt “For $q$, the generated answer $\Phi(q|a)$ matches the ground truth $R$ and is correct”. The reward is computed from the model’s logits, providing a scalar score that enables fine-grained, differentiable supervision over diverse reasoning trajectories.
With the reward $r$, the AdaReasoner is optimized using the gradient descent (REINFORCE) algorithm [43], where the overall policy $\Pi_{\Theta}(a\mid q,\Phi)$ is factorized into three heads with a shared feature extractor $f_{\theta_{c}}$, and $\Theta=\{\theta_{c},\theta_{p},\theta_{t},\theta_{s}\}$ denotes the complete set of trainable parameters. For each head $j\in\{p,t,s\}$, we define the head-specific loss as $\mathcal{L}_{j}=-r\log\Pi_{\theta_{j}}(a\mid q,\Phi)$, resulting in a total loss $\mathcal{L}=\sum_{j\in\{p,t,s\}}\mathcal{L}_{j}$. The gradients are then computed via the chain rule, where the shared-layer gradient is aggregated as $\nabla_{\theta_{c}}\mathcal{L}=\sum_{j\in\{p,t,s\}}\nabla_{\theta_{c}}\mathcal{L}_{j}$, and used for updating
@@
This training scheme ensures that each sub-policy is guided by its own loss while the shared feature extractor $f_{\theta_{c}}$ is jointly optimized by all heads, thereby promoting coherence across the three action dimensions and preventing convergence to conflicting optima, in line with findings from multi-task learning [38]. Further training details are described in Algorithm 1.